calculate the solubility of zns in 1 f acetic acid


1
Expert's answer
2021-03-30T07:48:25-0400

Molar Mass of ZnS = 97.474 g/mol

Molar Mass of CH3COOH = 60.052

1f = N /60.052

N = 60.052× 1 = 60.052

= 60.052 + 97.474 = 157.526

= (157.526/60.052) × 100g

= 262.32g/100g